OpenStudy (anonymous):

which is a possible solution of 6+ square root of 3x^4+1 =10-2x^2 ?

OpenStudy (anonymous):

subtract 6 from both sides and then square the result. solve for x

OpenStudy (anonymous):

\[6+\sqrt{3x^4+1}=10-2x^2\] \[\sqrt{3x^4+1}=4-2x^2\] \[3x^4+1=(4-2x^2)^2=4 x^4-16 x^2+16\]

OpenStudy (anonymous):

\[x^4-16x^2+15=0\] \[(x^2-1)(x^2-15)=0\] \[x^2=1\] \[x=\pm1\] \[x^2=15\] \[x=\pm\sqrt{15}\]

OpenStudy (anonymous):

so we have 4 "possible" solutions, but you have to check them to see which if any are correct

OpenStudy (anonymous):

i think just 1 and -1 work

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

yw